OpenStudy (anonymous):

i need to find the slope of (-4,-17) and (-14,-18) if it exists. I am just having a hard time with this..

OpenStudy (anonymous):

slope = 1/10

OpenStudy (anonymous):

slope is change in y over change in x, or (y2-y1)/(x2-x1). So (-17 - -18)/(-4 - -14) = 1/10

OpenStudy (anonymous):

(1/10)

OpenStudy (anonymous):

Thank you